# On-Orbit Servicing-Integrated Maintenance Strategy for Satellite Constellation

### v1

@@ 2.3 Parking Spares
$\rho_{\mathrm{parking}}=1-\frac{\mathrm{ES}_{\mathrm{parking}}}{k_{Q}}.$ (28)
@@ 2.4 Annual Maintenance Cost and Mean Time to Disposal
$T_{\mathrm{oos}}=\frac{N_{\mathrm{oos}}}{\mu_{\mathrm{oos}}}.$ (40)
2.5 Model Validation
To evaluate the accuracy of our inventory management model, we compared its parameter estimates to the mean of 100 Monte Carlo simulation results for 25 randomly selected instances. Each simulation spans 60 years and is discretized into one-day intervals. The fixed parameters and trade space for randomly sampled parameters are in Tables 2 and 3. We repeatedly sampled the varying parameters until we obtained 25 instances satisfying the following conditions:
$\rho_{\mathrm{plane}}\geq 0.98,$ (41)
$\rho_{\mathrm{parking}}\geq 0.98,$ (42)
$T_{d}/N_{t}\leq T_{\mathrm{life}},$ (43)
$s\leq Q,$ (44)
$k_{s}\leq k_{Q}.$ (45)
Here, Eqs. (41) and (42) represent the rare shortage assumptions in Section 2.1; Eq. (43) imposes the lifespan condition; and Eqs. (44) and (45) maintain the inventory’s cyclic behavior.
@@ 2.5 Model Validation
Fuel mass conversion coefficient, M\$/kg $\epsilon_{\mathrm{fuel}}$ 0.01
Satellite production cost, M\$/satellite $c_{\mathrm{sat}}$ 0.5
Spare/failed satellite holding cost, M\$/satellite/year $c_{\mathrm{hold}}$ 0.5
Launch cost, M\$/launch $c_{\mathrm{lau}}$ 67
@@
Number of orbital planes in the constellation, planes $h_{\mathrm{plane}}$ $\left[20,40\right]$
@@
The absolute error evaluates rational parameters (e.g., $\rho_{\mathrm{plane}}$ and $\rho_{\mathrm{parking}}$):
@@
The results in Table 4 indicate that relative errors remain below 2% and absolute errors below 1%p, meaning that the constructed model accurately estimates the parameters.
Table 4: Averaged errors of result parameters
Result Notation Mean Error
@@
Mean stock level at in-plane spare $\overline{\mathrm{SL}}_{\mathrm{plane}}$ 1.95
Mean stock level at parking spare $\overline{\mathrm{SL}}_{\mathrm{parking}}$ 0.74
Mean stock level of failed satellites waiting for service $\overline{\mathrm{SL}}_{\mathrm{wait}}$ 1.85
Order frequency at in-plane spare $n_{\mathrm{plane}}$ 0.49
Order frequency at parking spare $n_{\mathrm{parking}}$ 1.88
Service frequency $n_{\mathrm{oos}}$ 0.66
Mean time to disposal of satellites serviced up to the maximum $T_{d}$ 0.35
Annual maintenance cost $\mathrm{AMC}$ 0.71
@@
Order fill rate at in-plane spare $\rho_{\mathrm{plane}}$ 0.04
Order fill rate at parking spare $\rho_{\mathrm{parking}}$ 0.05
@@ 3 Decision-Making Scenarios and Optimization Problem Formulations
Based on the model, we develop decision-making scenarios involving two stakeholders—the constellation operator and the OOS provider—and discuss potential tension between them. This section presents the details of the decision-making contexts under consideration and their mathematical formulations as optimization problems.
3.1 Constellation Operator
The goal of the constellation operator is to maintain their system at an appropriate service level while minimizing relevant costs. Such a decision-making context can be formulated as an optimization problem, $\mathbf{P}_{1}$.
($\mathbf{P}_{1}$) Minimizing the Annual Maintenance Cost of the Constellation
$\min\ \mathrm{AMC}\left(\mathbf{x}_{\mathrm{con}};\mathbf{q}_{1}\right),$ (48)
@@ 3.1 Constellation Operator
$\rho_{\mathrm{plane}}\left(\mathbf{x}_{\mathrm{con}};\mathbf{q}_{1}\right)\geq\rho_{\mathrm{plane}}^{\mathrm{req}}$ (49)
$\rho_{\mathrm{parking}}\left(\mathbf{x}_{\mathrm{con}};\mathbf{q}_{1}\right)\geq\rho_{\mathrm{parking}}^{\mathrm{req}}$ (50)
@@
Eqs. (43)-(45),
$\mathbf{x}_{\mathrm{con}}=\left[s,Q,k_{s},k_{Q},N_{\mathrm{oos}},N_{\mathrm{parking}},h_{\mathrm{parking}}\right]\in\mathbb{Z}_{++}^{5}\times\mathcal{H}_{\mathrm{parking}}.$ (52)
Here, the vector of decision variables, $\mathbf{x}_{\mathrm{con}}$ in Eq. (52), includes parameters related to the replenishment policy and the configuration of parking orbits, while $\mathbf{q}_{1}$ contains the remaining parameters. Equation (48) defines the objective of the problem (minimizing the annual system maintenance cost). Equations (49) and (50) ensure that the maintenance system maintains an appropriate service level. Equation (51) constrains the number of loaded satellites so that it does not exceed the launch capacity ($Q_{\max}$). Finally, $\mathcal{H}_{\mathrm{parking}}$ denotes the set of available parking orbit altitudes.
3.2 OOS Provider
The main goal of an OOS provider is to maximize profit. The annual profits (AP) of the OOS provider are defined as
@@ 3.2 OOS Provider
where $p_{\mathrm{oos}}$ and $c_{\mathrm{oos}}$ are the unit price and cost of an OOS service, and $n_{\mathrm{oos}}N_{\mathrm{plane}}$ corresponds to the total annual number of services performed.
Given the OOS technologies, service types, and target subsystems of the spacecraft to be serviced, the OOS provider’s operational effort determines the service’s responsiveness and cost. For example, if the OOS provider allocates more resources (e.g., additional servicing spacecraft, depots, and propellant for faster orbital transfers), the servicing cost increases, but so does responsiveness. While a detailed investigation of operational rule-setting and cost estimation should be a separate research topic, here we introduce a first-cut model representing the relationship between cost and responsiveness as follows:
@@
Here, $c_{\mathrm{oos},\min}$ (>0>0) denotes the lower bound of the cost, $\mu_{\mathrm{oos},\mathrm{ideal}}$ (>0>0) represents the upper bound of responsiveness, and $\alpha_{1}$ (>0>0) and $\alpha_{2}$ (>0>0) are shape parameters. This function is strictly decreasing with respect to MTTR, exhibiting knee points and two asymptotes parallel to the cost and MTTR axes. Figure 3 illustrates the overall trend and the effect of each parameter.
Such a relationship, where increased effort yields greater outcomes but with a diminishing rate of increment, reflects the trade-off between service cost and responsiveness. This behavior is consistent with empirical observations in terrestrial logistics [30] and prior studies on OOS [32, 31, 33].
@@
With the aid of the cost-responsiveness function, we formulate the optimization problem $\mathbf{P}_{2}$, representing the decision-making of the OOS provider:
($\mathbf{P}_{2}$) Maximizing Profits of OOS Operations
$\max\ \mathrm{AP}\left(\mathbf{x}_{\mathrm{oos}};\mathbf{q}_{2}\right),$ (55)
subject to
$c_{oos}\left(\mu_{\mathrm{oos}};\mathbf{q}_{2}\right)\leq p_{\mathrm{oos}},$ (56)
$\mathrm{AMC}\left(\mathbf{x}_{\mathrm{oos}};\mathbf{q}_{2}\right)\leq\mathrm{AMC}_{\mathrm{ref}},$ (57)
$\mathbf{x}_{\mathrm{oos}}=\left[p_{\mathrm{oos}},\mu_{\mathrm{oos}}\right]\in\mathbb{R}_{++}^{2}.$ (58)
Here, the vector of decision variables $\mathbf{x}_{\mathrm{oos}}$ in Eq. (58) consists of two components: the price of the OOS per service, $p_{\mathrm{oos}}$, and the service responsiveness, $\mu_{\mathrm{oos}}$. Note that $\mu_{\mathrm{oos}}$ is the inverse of the MTTR. The vector $\mathbf{q}_{2}$ contains the remaining parameters, including those associated with the maintenance strategy of the constellation operator, which served as decision variables in $\mathbf{P}_{1}$. The objective function in Eq.(55) seeks to maximize the profits. Equation (56) ensures that the OOS price is no less than the corresponding service cost. Equation (56) ensures that the OOS price is no less than the corresponding service cost. Equation (57) constrains the OOS provider to set its service charge and responsiveness within a range that prevents the constellation operator from abandoning the OOS-integrated strategy. The term $\mathrm{AMC}{\mathrm{ref}}$ serves as a reference maintenance cost for adopting OOS; it can be interpreted as the maintenance cost without OOS or as a value that includes a premium over the reference cost from an alternative, non-OOS strategy. This premium reflects the additional complexity and uncertainty introduced by OOS integration, for which the operator may require greater expected benefits.
3.3 Tension Between Constellation Operator & OOS Provider
Previous subsections introduced the optimization problems reflecting the decision-making contexts of two main stakeholders: the constellation operator and the OOS provider. Note that these two problems are highly interdependent. On one hand, if the OOS provider raises the OOS price to increase profitability, the constellation operator would rely less on OOS for system maintenance and instead deploy new satellites. On the other hand, if the constellation operator relies more on OOS, the profitability of the OOS provider would improve without changing its decisions. Balancing such tensions and designing appropriate incentives to achieve mutually beneficial outcomes that stimulate the overall space industry are crucial challenges. Indeed, the interdependence between satellite operators and OOS providers is sometimes regarded as a chicken-and-egg problem [34].
The following bi-objective optimization problem expresses the tension between the constellation operator and the OOS provider by combining $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$:
($\mathbf{P}_{3}$) Bi-Objective Optimization for Multiple Stakeholders
$\min\ \mathrm{AMC}\left(\mathbf{x};\mathbf{q}_{3}\right),$ (59)
$\max\ \mathrm{AP}\left(\mathbf{x};\mathbf{q}_{3}\right),$ (60)
subject to
$Qk_{Q}\leq Q_{\max},$ (61)
$c_{\mathrm{oos}}\left(\mu_{\mathrm{oos}};\mathbf{q}_{3}\right)\leq p_{\mathrm{oos}},$ (62)
$\mathrm{AMC}\left(\mathbf{x};\mathbf{q}_{3}\right)\leq\mathrm{AMC}_{\mathrm{ref}},$ (63)
Eqs. (43)-(45),
$\mathbf{x}=\left[\mathbf{x}_{\mathrm{con}},\mathbf{x}_{\mathrm{oos}}\right]\in\mathbb{Z}_{++}^{5}\times\mathcal{H}_{\mathrm{parking}}\times\mathbb{R}_{++}^{2}.$ (64)
Here, the vector of decision variables, $\mathbf{x}$ in Eq. (64), is the concatenation of $\mathbf{x}_{\mathrm{con}}$ and $\mathbf{x}_{\mathrm{oos}}$ from $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$, respectively. Solving this bi-objective optimization problem yields a family of non-dominated solutions—solutions that are not inferior to any other solution with respect to all objective functions [35]—known as the Pareto front [36, 37]. To obtain the Pareto front, we employ the non-dominated sorting genetic algorithm II (NSGA-II), a metaheuristic based on the genetic algorithm [38].

### v2

@@ 2.4 Parking Spares
$\rho_{\mathrm{parking}}=1-\frac{\mathrm{ES}_{\mathrm{parking}}}{k_{Q}}{\color[rgb]{0,0,0}{,}}$ (28)
where $N_{t}$ is the number of discretized time units per year.
@@ 2.5 Annual Maintenance Cost and Mean Time to Disposal
$T_{\mathrm{oos}}=\frac{N_{\mathrm{oos}}}{\mu_{\mathrm{oos}}}$ (40)
Note that $T_{\mathrm{parking}}$, $T_{\mathrm{plane}}$, and $T_{\mathrm{oos}}$ are obtained as the ratios of the corresponding stock levels to their outflow rates, $T_{\mathrm{trans}}$ follows from the definition of the mean of a continuous random variable with probability density function $f_{\mathrm{plane}}$, and $T_{\mathrm{opr}}$ is given by the reciprocal of the satellite failure rate. The factors ($N_{\mathrm{oos}}+1$) and $N_{\mathrm{oos}}$ arise because the satellite passes through the in-plane spare and operational phases ($N_{\mathrm{oos}}+1$) times, whereas it undergoes on-orbit servicing only $N_{\mathrm{oos}}$ times before disposal.
3 Model Validation
The analytical model for the performance parameters constructed in Section 2 is validated by comparing its values with those obtained from discrete simulations. We first describe the generation of the validation instances and then present the results.
3.1 Instance Generation
To evaluate the accuracy of the proposed inventory management model, we compared its estimates with the sample means obtained from $100$ Monte Carlo simulation runs. Each simulation spans $60$ years and is discretized into one-day intervals.
Some parameters were held fixed, while others were randomly sampled. Table 2 lists the fixed parameters, and Table 3 gives the ranges for the sampled parameters. Most parameter values were adopted from previous work [32], whereas the OOS-related parameters were defined in this study to span a broad range of plausible values. The OOS price, $p_{\mathrm{oos}}$, was set to \$22M, which is approximately equal to the sum of the satellite production cost, $c_{\mathrm{sat}}=$\$0.50.5M, and the per-satellite launch cost assuming a launch capacity of $40$ satellites, i.e., $c_{\mathrm{lau}}/40\approx$\$1.71.7M. The same launch capacity was adopted as the reference launcher specification in the case study, following previous work [32]. The serviceable failure fraction, $r_{\mathrm{oos}}$, was varied from $0.1$ to $0.9$. The OOS MTTR, $\mu_{\mathrm{oos}}^{-1}$, was varied from $1$ week to $3$ months to represent a wide range of plausible service-performance levels in LEO [77]. The maximum number of services, $N_{\mathrm{oos}}$, was varied from $1$ to $4$, based on the feasible range under the fixed satellite lifetime of $T_{\mathrm{life}}=30$ years.
@@ 3.1 Instance Generation
Fuel mass conversion coefficient, \$M/kg $\epsilon_{\mathrm{fuel}}$ 0.01
Satellite production cost, \$M/satellite $c_{\mathrm{sat}}$ 0.5
Spare/failed satellite holding cost, \$M/satellite/year $c_{\mathrm{hold}}$ 0.5
Launch cost, \$M/launch $c_{\mathrm{lau}}$ 67
@@
Number of orbital planes in the constellation, planes $N_{\mathrm{plane}}$ $\left[20,40\right]$
@@
We repeatedly sampled the varying parameters until we obtained $100$ instances satisfying the following conditions:
$\rho_{\mathrm{plane}}\geq\rho_{\mathrm{plane}}^{\mathrm{req}},$ (41)
$\rho_{\mathrm{parking}}\geq\rho_{\mathrm{parking}}^{\mathrm{req}},$ (42)
$T_{d}/N_{t}\leq T_{\mathrm{life}},$ (43)
$s\leq Q,$ (44)
$k_{s}\leq k_{Q}.$ (45)
Here, Eqs. (41) and (42) represent the rare shortage assumption (Assumption 7) in Section 2.2 and are also used for service level requirements in later sections. Equation (43) imposes the lifespan condition, and Eqs. (44) and (45) maintain the cyclic behavior of the inventory. For $\rho_{\mathrm{plane}}^{\mathrm{req}}$ and $\rho_{\mathrm{parking}}^{\mathrm{req}}$, we considered values of $0.98$, $0.95$, and $0.90$, which yielded three sets of $100$ instances for testing stress cases under the rare shortage assumption.
3.2 Results
We tracked the mean stock levels of in-plane spares ($\overline{\mathrm{SL}}_{\mathrm{plane}}$), parking spares ($\overline{\mathrm{SL}}_{\mathrm{parking}}$), and failed satellites awaiting service ($\overline{\mathrm{SL}}_{\mathrm{wait}}$); the order frequencies at the in-plane and parking spare echelons ($n_{\mathrm{plane}}$ and $n_{\mathrm{parking}}$); the OOS frequency ($n_{\mathrm{oos}}$); the mean time to disposal for satellites serviced up to the maximum allowable number of times ($T_{d}$); the annual maintenance cost ($\mathrm{AMC}$); and the order fill rates of the in-plane and parking spare echelons ($\rho_{\mathrm{plane}}$ and $\rho_{\mathrm{parking}}$).
For the rate parameters, $\rho_{\mathrm{plane}}$ and $\rho_{\mathrm{parking}}$, we used the absolute error:
@@ 3.2 Results
To construct stress cases for Assumption 9, which assumes an exponential distribution for the OOS lead time, we performed simulations using alternative distributions for lead time with the same MTTR, $\mu_{\mathrm{oos}}^{-1}$, but different coefficients of variation (CVs). Here, CV is defined as the ratio of the standard deviation to the mean. The tested cases were deterministic, gamma, and lognormal. For the gamma distribution, we considered CV values of $1/2$, $\sqrt{2}, 11$, $\sqrt{2}$, and $2$, whereas for the lognormal distribution, we considered CV values of $\sqrt{2}$, $2$, and $4$. The deterministic case corresponds to a fixed lead time equal to $\mu_{\mathrm{oos}}^{-1}$. The gamma distribution was included because it generalizes the exponential distribution, with CV =1=1 corresponding to the baseline exponential case (Assumption 9). Therefore, both lower- and higher-variability cases relative to this baseline were tested. The lognormal distribution was included to examine long-tail stress cases, with CV values exceeding the baseline value of $1$. The results are summarized in Tables 46, corresponding to the cases where $\rho_{\mathrm{plane}}^{\mathrm{req}}=\rho_{\mathrm{parking}}^{\mathrm{req}}=0.98$, $0.95$, and $0.90$, respectively.
Table 4: Averaged errors of result parameters, $\rho_{\mathrm{plane}}^{\mathrm{req}}=\rho_{\mathrm{parking}}^{\mathrm{req}}=0.98$
Deterministic Gamma Lognormal
$\mathrm{CV}$ 0 $1/2$ $\sqrt{2} 11$ (Exponential) $\sqrt{2}$ 22 $\sqrt{2}$ 22 44
@@
$\overline{\mathrm{SL}}_{\mathrm{plane}}$ 1.89 1.93 1.92 1.90 1.96 1.91 1.89 1.86 1.87
$\overline{\mathrm{SL}}_{\mathrm{parking}}$ 0.73 0.76 0.75 0.66 0.76 0.70 0.75 0.76 0.82
$\overline{\mathrm{SL}}_{\mathrm{wait}}$ 1.08 1.95 2.01 2.06 2.51 2.75 2.01 2.29 3.04
$n_{\mathrm{plane}}$ 0.54 0.50 0.52 0.56 0.45 0.47 0.55 0.54 0.50
$n_{\mathrm{parking}}$ 1.51 1.71 1.64 1.40 1.46 1.38 1.54 1.72 1.53
$n_{\mathrm{oos}}$ 0.64 0.76 0.75 0.70 0.76 0.70 0.66 0.77 0.67
$T_{d}$ 0.43 0.42 0.46 0.39 0.44 0.46 0.48 0.48 0.48
$\mathrm{AMC}$ 0.62 0.60 0.60 0.60 0.66 0.59 0.56 0.66 0.70
@@
$\rho_{\mathrm{plane}}$ 0.10 0.10 0.10 0.09 0.10 0.11 0.10 0.11 0.10
$\rho_{\mathrm{parking}}$ 0.07 0.06 0.05 0.05 0.06 0.07 0.07 0.06 0.06
Table 5: Averaged errors of result parameters, $\rho_{\mathrm{plane}}^{\mathrm{req}}=\rho_{\mathrm{parking}}^{\mathrm{req}}=0.95$
Deterministic Gamma Lognormal
$\mathrm{CV}$ 0 $1/2$ $\sqrt{2} 11$ (Exponential) $\sqrt{2}$ 22 $\sqrt{2}$ 22 44
Relative Errors (%)
$\overline{\mathrm{SL}}_{\mathrm{plane}}$ 2.65 2.71 2.71 2.65 2.72 2.72 2.69 2.64 2.67
$\overline{\mathrm{SL}}_{\mathrm{parking}}$ 0.84 0.82 0.85 0.81 0.88 0.76 0.87 0.80 0.89
$\overline{\mathrm{SL}}_{\mathrm{wait}}$ 1.18 1.83 1.92 2.03 2.35 2.53 1.98 2.21 2.95
$n_{\mathrm{plane}}$ 0.55 0.51 0.49 0.54 0.44 0.47 0.56 0.51 0.47
$n_{\mathrm{parking}}$ 1.38 1.51 1.52 1.33 1.39 1.37 1.38 1.63 1.38
$n_{\mathrm{oos}}$ 0.69 0.78 0.75 0.70 0.77 0.75 0.66 0.77 0.66
$T_{d}$ 0.44 0.44 0.46 0.42 0.46 0.44 0.47 0.42 0.45
$\mathrm{AMC}$ 0.71 0.64 0.69 0.63 0.72 0.63 0.64 0.69 0.71
Absolute Errors (%p)
$\rho_{\mathrm{plane}}$ 0.34 0.35 0.36 0.33 0.35 0.35 0.35 0.34 0.32
$\rho_{\mathrm{parking}}$ 0.10 0.10 0.09 0.10 0.09 0.08 0.10 0.09 0.11
Table 6: Averaged errors of result parameters, $\rho_{\mathrm{plane}}^{\mathrm{req}}=\rho_{\mathrm{parking}}^{\mathrm{req}}=0.90$
Deterministic Gamma Lognormal
$\mathrm{CV}$ 0 $1/2$ $\sqrt{2} 11$ (Exponential) $\sqrt{2}$ 22 $\sqrt{2}$ 22 44
Relative Errors (%)
$\overline{\mathrm{SL}}_{\mathrm{plane}}$ 2.74 2.83 2.83 2.73 2.78 2.81 2.80 2.72 2.74
$\overline{\mathrm{SL}}_{\mathrm{parking}}$ 0.92 0.83 0.93 0.85 0.97 0.86 0.87 0.84 0.93
$\overline{\mathrm{SL}}_{\mathrm{wait}}$ 1.17 1.95 2.13 2.21 2.57 2.74 2.08 2.30 3.19
$n_{\mathrm{plane}}$ 0.55 0.51 0.56 0.53 0.49 0.52 0.59 0.56 0.53
$n_{\mathrm{parking}}$ 1.42 1.44 1.57 1.32 1.40 1.27 1.25 1.61 1.29
$n_{\mathrm{oos}}$ 0.73 0.82 0.79 0.70 0.82 0.76 0.71 0.80 0.72
$T_{d}$ 0.45 0.46 0.48 0.43 0.49 0.49 0.48 0.45 0.48
$\mathrm{AMC}$ 0.74 0.67 0.70 0.65 0.71 0.66 0.67 0.71 0.72
Absolute Errors (%p)
$\rho_{\mathrm{plane}}$ 0.43 0.44 0.44 0.40 0.42 0.42 0.45 0.41 0.39
$\rho_{\mathrm{parking}}$ 0.16 0.13 0.14 0.12 0.12 0.13 0.15 0.13 0.13
The results show that the proposed model remains a good approximation across all tested cases. In the baseline case corresponding to Assumption 9, i.e., the gamma distribution with CV =1=1 (equivalently, the exponential distribution), all reported relative errors remain below 33% for the three service level requirements, and both absolute order fill rate errors remain below 0.50.5%p. This indicates that the model reproduces the simulation results well under the nominal assumption.
Among the reported parameters, $\overline{\mathrm{SL}}_{\mathrm{plane}}$ and $\overline{\mathrm{SL}}_{\mathrm{wait}}$ exhibit the largest errors. This is expected because these quantities are most directly coupled to the OOS-enabled closed-loop recovery process. However, their sensitivities to lead-time variability differ. In particular, $\overline{\mathrm{SL}}_{\mathrm{wait}}$ is the most sensitive to the shape of the lead-time distribution: for both the gamma and lognormal families, its error increases as the CV increases, even though the MTTR is held fixed. This behavior is consistent with the fact that $\overline{\mathrm{SL}}_{\mathrm{wait}}$ depends most directly on the residence time in the OOS stage. By contrast, the error in $\overline{\mathrm{SL}}_{\mathrm{plane}}$ is relatively insensitive to the CV because the in-plane spare stock is influenced not only by recovery through OOS but also by replenishment from parking spares.
As the required order fill rate is relaxed from $0.98$ to $0.90$, the overall error level increases modestly, but the deterioration remains limited. Even in the most severe stress case—a lognormal lead time distribution with CV =4=4 under $\rho_{\mathrm{plane}}^{\mathrm{req}}=\rho_{\mathrm{parking}}^{\mathrm{req}}=0.90$—the largest average relative error is only 3.193.19%, observed for $\overline{\mathrm{SL}}_{\mathrm{wait}}$, while the largest average absolute error in order fill rate across all tested cases is only 0.450.45%p. Therefore, although higher-order features of the lead time distribution do affect the prediction of the waiting stock, the resulting loss of accuracy remains small for the tested service-level requirements, including thresholds as low as $0.90$.
In summary, these results support the use of Assumption 9 as a tractable approximation for system-level modeling. Model accuracy is driven primarily by the mean OOS lead time, whereas the exact distributional form plays a secondary role. Accordingly, the exponential assumption provides a reasonable balance between analytical simplicity and predictive accuracy for design-space exploration and managerial trade-off analysis, although some caution is warranted when interpreting $\overline{\mathrm{SL}}_{\mathrm{wait}}$ under highly variable, long-tailed lead-time distributions and low service-level requirements.
@@ 4 Decision-Making Scenarios and Optimization Problem Formulations
This study considers two primary stakeholders: the constellation operator and the OOS provider. The constellation operator aims to maintain the required service level of the constellation while minimizing maintenance cost. The OOS provider, in turn, aims to maximize profits by selecting service price and performance levels—in this study, responsiveness—that remain commercially attractive to the constellation operator. The decisions of the two stakeholders are strongly coupled: the operator’s maintenance strategy determines the demand for OOS, whereas the provider’s pricing and responsiveness influence the operator’s preferred maintenance strategy. Balancing these tensions and designing appropriate incentives to achieve mutually beneficial outcomes that support the broader space industry are important challenges. Indeed, the interdependence between satellite operators and OOS providers is sometimes regarded as a chicken-and-egg problem [78].
We first represent this interaction through the following bi-objective optimization problem, whose components are detailed in this section:
($\mathbf{P}$) Bi-Objective Optimization for the Constellation Operator and OOS Provider
$\min\ \mathrm{AMC}\left(\mathbf{x};\mathbf{q}\right),$ (48)
$\max\ \mathrm{AP}\left(\mathbf{x};\mathbf{q}\right),$ (49)
@@
Eqs. (41)–(45),
@@
$c_{\mathrm{oos}}\left(\mu_{\mathrm{oos}};\mathbf{q}\right)\leq p_{\mathrm{oos}},$ (51)
$\mathrm{AMC}\left(\mathbf{x};\mathbf{q}\right)\leq\mathrm{AMC}_{\mathrm{ref}},$ (52)
$\mathbf{x}=\left[s,Q,k_{s},k_{Q},N_{\mathrm{oos}},N_{\mathrm{parking}},h_{\mathrm{parking}},p_{\mathrm{oos}},\mu_{\mathrm{oos}}\right]\in\mathbb{Z}_{++}^{6}\times\mathcal{H}_{\mathrm{parking}}\times\mathbb{R}_{++}^{2}.$ (53)
Here, $\mathbf{x}$ denotes the unified decision vector, and $\mathbf{q}$ collects all remaining parameters. The decision vector contains both the constellation operator’s decision variables, $\left[s,Q,k_{s},k_{Q},N_{\mathrm{oos}},N_{\mathrm{parking}},h_{\mathrm{parking}}\right]$, and the OOS provider’s decision variables, $\left[p_{\mathrm{oos}},\mu_{\mathrm{oos}}\right]$. In addition, $\mathcal{H}_{\mathrm{parking}}$ denotes the set of available parking orbit altitudes.
From the constellation operator’s perspective, Eq. (48) minimizes the annual maintenance cost. The service-level requirements in Eqs. (41) and (42) ensure that the maintenance system achieves the required order fill rates for in-plane and parking spares, respectively. Equation (50) limits the number of loaded satellites so that it does not exceed the launch capacity, $Q_{\max}$, while Eqs. (43)–(45) impose the validity conditions of the analytical model.
From the OOS provider’s perspective, Eq. (49) represents the maximization of annual profit, defined as
@@
where $p_{\mathrm{oos}}$ and $c_{\mathrm{oos}}$ are the unit price and cost of an OOS, respectively. The term $n_{\mathrm{oos}}N_{\mathrm{plane}}$ represents the total annual number of services performed, computed as the product of the OOS frequency, $n_{\mathrm{oos}}$, and the number of target orbits, $N_{\mathrm{plane}}$. Thus, annual profit equals the profit per service multiplied by the total annual service volume.
Given the OOS technologies, service types, and target subsystems, higher responsiveness generally requires greater operational effort, such as additional servicers, depots, and propellant for faster orbital transfers. To capture this trade-off, we adopt the following first-cut model relating cost and responsiveness:
@@
Here, $c_{\mathrm{oos},\min}$ (>0>0) denotes the lower bound of the service cost, $\mu_{\mathrm{oos},\mathrm{ideal}}$ (>0>0) denotes the upper bound of service responsiveness, and $\alpha_{1}$ (>0>0) and $\alpha_{2}$ (>0>0) are shape parameters. This function is strictly decreasing with respect to MTTR and exhibits diminishing returns as responsiveness approaches its ideal bound. Such a relationship, in which greater operational effort yields improved responsiveness at a diminishing rate, is consistent with empirical observations in terrestrial logistics [79] and prior OOS-related studies [50, 77, 51].
@@
Equation (51) ensures that the OOS price is no less than the corresponding service cost. Equation (52) constrains the provider to choose the price and responsiveness so that the resulting OOS-integrated maintenance strategy remains acceptable to the constellation operator. The reference value $\mathrm{AMC}_{\mathrm{ref}}$ may be interpreted as the maintenance cost of a benchmark strategy without OOS, or as a premium-adjusted reference value reflecting the additional complexity and uncertainty introduced by OOS integration.
Solving Problem $\mathbf{P}$ yields a family of non-dominated solutions[80], known as the Pareto front [81, 82]. In this study, the Pareto front is obtained using the non-dominated sorting genetic algorithm II (NSGA-II), a genetic algorithm-based metaheuristic for multi-objective optimization [83].